Henderson Inc. plans to introduce a new product next year. This product has a two-year life and an estimated demand of 20,000 units annually. The product will be produced 50 weeks each year. Henderson, Inc. estimates the following costs:

• Direct materials will be $25 per unit.
• Setup costs will be $300 per set up, five setups will be required per week.
• Specialized equipment must be rented for $7,000 per week.
• Design costs are estimated to be $50,000
• Research and development costs are estimated at $350,000
• Labor will be paid at $15 per hour.

Henderson, Inc still has the same costs but now wants to the determine the selling price using Target pricing. They have set the price at $100 per unit and require a 60% return.

1. Will the product meet their expectations for a 60% return?
2. What if they required a 25% return?

Note that “Return” refers to the percentage a company wants left over from the selling price after the company has factored in what it costs to make it. in other words
Return = Selling price minus costs.
Answer #1

per unit cost =

Direct material = $25

Set up cost = 300*5*50/20000 = $ 3.75

equipment rent per unit = 7000*50/20000 = $17.5

Design cost per unit = 50000/20000 = $2.5

R& D cost = 350000 (Ignored as sunk cost in nature )

Labour cost = $15

Total cost per unit = $63.75

Required return = 60% = $38.25

Target price = cost price + desired return = 102

If they set price = 102 , desired return shall not be achieved.

2. if 25% return required then desired price shall be =63.75 + 63.75*25% = 79.69 i.e around $80